# Your Finetuned Large Language Model is Already a Powerful Out-of-distribution Detector

### v1

@@ Abstract
We revisit the likelihood ratio between a pretrained large language model (LLM) and its finetuned variant as a criterion for out-of-distribution (OOD) detection. The intuition behind such a criterion is that, the pretrained LLM has the prior knowledge about OOD data due to its large amount of training data, and once finetuned with the in-distribution data, the LLM has sufficient knowledge to distinguish their difference. Leveraging the power of LLMs, we show that, for the first time, the likelihood ratio can serve as an effective OOD detector. Moreover, we apply the proposed LLM-based likelihood ratio to detect OOD questions in question-answering (QA) systems, which can be used to improve the performance of specialized LLMs for general questions. Given that likelihood can be easily obtained by the loss functions within contemporary neural network frameworks, it is straightforward to implement this approach in practice. Since both the pretrained LLMs and its various finetuned models are available, our proposed criterion can be effortlessly incorporated for OOD detection without the need for further training. We conduct comprehensive evaluation across on multiple settings, including far OOD, near OOD, spam detection, and QA scenarios, to demonstrate the effectiveness of the method.
@@ 1 Introduction
In OOD detection (Hendrycks & Gimpel, 2016), the term “in-distribution” refers specifically to the distribution of the training data. In natural language processing, OOD detection has been studied in small and task-specific models for settings such as translation (Xiao et al., 2020) and question answering (Lyu et al., 2020). With the advancement of large and general models like large language models (LLMs), the scope of training data has significantly broadened, positioning these large models embodied with general knowledge and intelligence as “base models”. In the era of large foundation models (Bommasani et al., 2021), the prevalent training paradigm has shifted from end-to-end learning towards finetuning a pretrained base model. This shift calls for a revisiting of the definition of “in-distribution” and its relationship to the training distribution of base models.
There is usually no prior information about the OOD data in conventional OOD detection (Hendrycks & Gimpel, 2016). Bishop (1994) introduced the idea that OOD data could be viewed as data coming from a distinct OOD distribution, suggesting the use of the likelihood ratio between this OOD distribution and the in-distribution as a detection criterion. However, given the absence of information about OOD data, Zhang & Wischik (2022) introduced the concept of OODProxy, a conceptual framework where the OOD distribution is represented by a proxy, incorporating what is known or assumed about the OOD data. From this perspective, it is commonly recognized that many OOD detection methods leveraging likelihood ratios are essentially applying their unique OOD proxy distributions, each reflecting different assumptions or prior knowledge concerning the nature of the OOD data (Ren et al., 2019; Serrà et al., 2019; Schirrmeister et al., 2020; Zhang et al., 2021).
In our paper, we propose that pretrained base models can function as a repository of prior knowledge for OOD data relative to in-distribution data, effectively acting as OOD proxy distributions. Guided by this insight, we discover that the likelihood ratio between the base model and its finetuned counterpart serves as an effective criterion for detecting OOD data. Moreover, for LLM-based question-answering (QA) systems, the same likelihood ratio excels in detecting OOD questions. By identifying and rejecting these OOD questions, we can greatly enhance the robustness of current QA systems.
The convenience of obtaining likelihood from loss functions in current neural networks enables a simple and straightforward implementation of this method in practice. Moreover, it is worth mentioning that numerous practitioners likely already have both a pretrained and a finetuned LLM at their disposal. This setup inherently equips them with the capacity for OOD detection without necessitating additional training efforts.
@@ 2 Background and Preliminaries
OOD Detection. We start with an in-distribution dataset, denoted as $\mathcal{D}_{\text{in}}$, assuming that the data within $\mathcal{D}_{\text{in}}$ is sampled from an in-distribution probability distribution $p_{\text{in}}$. The objective of OOD detection is to determine whether a given input data $x$ originates from $p_{\text{in}}$. Hendrycks & Gimpel (2016) were the first to highlight the significance of this problem in deep learning era and introduced a practical benchmark for evaluation. This involves training a model on $\mathcal{D}_{\text{in}}^{\text{train}}$, the training subset of the in-distribution dataset, with the model providing a detection criterion $S$ for identifying OOD data. We then gather a dataset from domains different from $\mathcal{D}_{\text{in}}$, labeled as $\mathcal{D}_{\text{out}}$. The effectiveness of the criterion $S$ is assessed by applying it to data from $\mathcal{D}_{\text{in}}^{\text{test}}\cup\mathcal{D}_{\text{out}}^{\text{test}}$ and evaluating the performance using metrics such as AUROC, AUPR, and FPR95 (Yang et al., 2022). These metrics help determine how well $S$ can differentiate between the in-distribution dataset and the OOD dataset. A high performance across these metrics signifies a robust OOD detection capability.
@@
In Hendrycks & Gimpel (2016)’s foundational study, it is assumed that for the in-distribution dataset $\mathcal{D}_{\text{in}}$, each data point is accompanied by a classification label. Our approach diverges slightly from this premise, as we operate under the assumption that no labels are available. This scenario is often referred to as “unsupervised OOD detection” or “OOD detection without in-distribution labels”.
@@
The Paradox in Unsupervised OOD Detection. In the context of unsupervised OOD detection, Nalisnick et al. (2018) revisit Bishop (1994)’s suggestion that a probabilistic generative model $p_{\theta}$ could model the in-distribution $p_{\text{in}}$, proposing to use the model output $S(x)=p_{\theta}(x)$ as a criterion for evaluating a given input $x$. Surprisingly, Nalisnick et al. discovered that in certain scenario - such as when $\mathcal{D}_{\text{in}}$ is CIFAR10 and $\mathcal{D}_{\text{out}}$ is SVHN, or $\mathcal{D}_{\text{in}}$ is FashionMNIST and $\mathcal{D}_{\text{out}}$ is MNIST - the OOD data received higher $p_{\theta}(x)$ scores than the in-distribution data. This counterintuitive finding has been labeled as a “paradox”.
@@
OOD Proxy. To address the paradox, the studies (Ren et al., 2019; Serrà et al., 2019; Schirrmeister et al., 2020; Zhang et al., 2021) have put forward the idea of utilizing the likelihood ratio as the criterion for identifying OOD data. Zhang & Wischik (2022) integrates these techniques into a comprehensive structure termed the OOD proxy framework. Within this framework, it is posited that in-distribution data can be characterized as samples from a distribution $p_{\text{in}}$, whereas OOD data are samples from a distribution $p_{\text{out}}$. According to the Neyman-Pearson lemma, the likelihood ratio represents the optimal criterion for OOD detection in theory, which is mathematically expressed as:
@@
where obtaining $p_{\text{out}}$ is challenging. To address this, Zhang & Wischik (2022) introduced the concept of utilizing a proxy distribution, $p_{\text{out}}^{\text{proxy}}$, which incorporates human subjective understanding of the OOD distribution. For instance, Ren et al. (2019) define $p_{\text{out}}^{\text{proxy}}$ as the distribution representing background statistics; Serrà et al. (2019) consider it as the distribution of data compression; Schirrmeister et al. (2020) describe $p_{\text{out}}^{\text{proxy}}$ as a general distribution; and for Zhang et al. (2021), $p_{\text{out}}^{\text{proxy}}$ corresponds to the distribution of a local autoregressive model.
In the OOD proxy framework, the use of likelihood as a detection criterion, following the approach of Bishop (1994) and Nalisnick et al. (2018), essentially assumes a uniform distribution for $p_{\text{out}}^{\text{proxy}}$. The suboptimal performance associated with this method is not surprising; it highlights the limitations of an improper prior assumption.
@@
Autoregressive language models (Brown et al., 2020) are types of probabilistic models that compute the likelihood of a sentence $x=x_{1},\dots,x_{T}$, where $T$ denotes the sentence length and $x_{t}$ represents each word at position $t$. By definition of conditional probability, the likelihood $p(x)=p(x_{1},\dots,x_{T})$ can be calculated as the product of conditional probabilities: $p(x_{1})p(x_{2}|x_{1})p(x_{3}|x_{1},x_{2})\dots p(x_{T}|x_{1},\dots,x_{T-1})$. Each term $p(x_{t}|x_{<t})$ represents the probability of predicting the subsequent word $x_{t}$ given all the previous words $x_{1},\dots,x_{t-1}$. In neural language models, these conditional probabilities are typically modeled using a softmax function over the output vocabulary.
@@
3 Pretrained Large Language Model as a OOD Proxy
@@
This criterion essentially employs the pretrained large language model as an OOD proxy introduced in Section 2. This strategy is particularly practical given the widespread availability of pretrained LLMs. Finetuning these models to adapt their distribution for specific domain contexts is a standard practice, meaning many practitioners may already possess finetuned LLMs that represent the distribution of their specific datasets. With both pretrained and finetuned models at hand, calculating the likelihood ratio becomes straightforward, eliminating the need for additional training. For instance, suppose we possess a LLM that has been finetuned on legal documents. Given a new document $x$, we can determine whether it is a legal document by utilizing the likelihood ratio $S(x)$.
@@
Utilizing a pretrained LLM as an OOD proxy stems from the rationale that the OOD proxy distribution should encapsulate general characteristics of OOD data, including prior knowledge or subjective insights about the OOD data. While it’s conceivable to presume OOD data emanate from a uniform distribution devoid of any prior knowledge, evidence presented by Nalisnick et al. (2018) indicates that relying on a uniform distribution as the OOD proxy can be ineffective in practice. In the context of language models, the well-known infinite monkey theorem [^1] The infinite monkey theorem states that a monkey randomly typing on a typewriter for an infinitely long period will almost surely produce any given text, such as the complete works of William Shakespeare. (Borel, 1913; Eddington, 2019) suggests that any text could theoretically be generated from a uniform distribution; however, this does not serve as an effective representation of any coherent language. Figure 1 depicts the relationship among sentences within a specific domain, the comprehensive set of human language, and all conceivable character permutations. In fact, coherent human language forms a minor subset of all potential character arrangements. Consequently, the assumption that OOD data ought to represent meaningful human language constitutes a strong prior.
Therefore, we advocate for the use of a pretrained LLM as a more suitable OOD proxy. Given their extensive parameters and training on vast corpora (for example, the Llama-2 model, as mentioned by Touvron et al. (2023), is trained on 7 trillion tokens), it is plausible to consider that an LLM encompasses the breadth of human language. Assuming the LLM estimates the distribution of all human language, it is logical to designate the pretrained LLM as the OOD proxy.
@@ 4 Likelihood Ratio OOD Detection for QA Systems
$S_{q}(q,a)=\frac{p_{\theta}(q)}{p_{\theta^{\prime}}(q)},\quad S_{a}(q,a)=\frac{p_{\theta}(a)}{p_{\theta^{\prime}}(a)},\quad S_{q,a}(q,a)=\frac{p_{\theta}(q,a)}{p_{\theta^{\prime}}(q,a)},$
$S_{a|q}(q,a)=\frac{p_{\theta}(a|q)}{p_{\theta^{\prime}}(a|q)}=\frac{p_{\theta}(q,a)p_{\theta^{\prime}}(q)}{p_{\theta^{\prime}}(q,a)p_{\theta}(q)}=\frac{S_{q,a}(q,a)}{S_{q}(q,a)},$
where $S_{q}$, $S_{a}$, $S_{q,a}$, and $S_{a|q}$ are defined as follows: $S_{q}$ is the likelihood ratio for the question, $S_{a}$ for the answer, $S_{q,a}$ for the question-answer pair, and $S_{a|q}$ for the answer given the question. All these criterion are the ratios between the likelihoods assigned by the finetuned model to those assigned by the base model.
Refer to captionRefer to captionRefer to caption
Figure 2: Example question-answer sets produced by MetaMath-7B. The responses to In-D questions are accurate and logical. However, for OOD questions, MetaMath-7B generates unreasonable answers, responding to a straightforward query with unnecessary mathematical calculations or producing repetitive sentences with no useful information.
5 Experiments and Results
We adhere to the definitions of far OOD and near OOD as outlined by Yang et al. (2022) in their work. Near OOD datasets exhibit only a semantic shift from the In-D datasets, whereas far OOD also encompasses a significant covariate (domain) shift. For far OOD evaluations, we designate two distinct datasets as In-D and OOD. For near OOD, we divide a single dataset into two groups: one serving as the In-D with certain classes and the other as OOD with a different set of classes.
Additionally, we demonstrate the capability of our proposed method in detecting OOD instances within the context of spam detection (Labonne & Moran, 2023) - a practical application for our unsupervised OOD detection technique, especially where In-D labels are absent. We show that our method achieves commendable results in spam detection even without access to any spam data. Moreover, when spam data is available and we further finetune the OOD proxy distribution using this data, our results are competitive with state-of-the-art (SOTA) spam detection algorithms.
Finally, we evaluate our approach within a real question-answering (QA) context, utilizing MetaMath - a Llama-2 model finetuned for math problem-solving, as described by Yu et al. (2023). By implementing the likelihood-ratio-based criteria outlined in Section 4, we find that for specific short questions, having the LLM provide an answer and subsequently applying a criterion that analyzes both the question and answer leads to consistently improved outcomes in identifying OOD questions.
Evaluation Metrics
We employ AUROC (Area Under the Receiver Operating Characteristic curve), AUPR (Area Under the Precision-Recall curve), and FPR95 (False Positive Rate at 95% True Positive Rate) as our evaluation metrics. These metrics are commonly utilized in evaluating the performance of OOD detection methods (Hendrycks & Gimpel, 2016; Yang et al., 2022).
5.1 Far OOD Detection
For text data, detecting far OOD instances is relatively less challenging. Uppaal et al. (2023) have illustrated that utilizing the latent distance from a pretrained RoBERTa model significantly addresses far OOD detection, especially when the 20 Newsgroups (20NG) dataset serves as the in-distribution. In this context, we present that our proposed method also attains nearly perfect performance under the same experimental conditions, as detailed in Table 1.
Note that, the notation ‘Model-XB LH/LR’ is used, where ‘Model’ can be either Llama (Touvron et al., 2023) or Mistral (Jiang et al., 2023), denoting two popular open-source large language models (LLMs). Here, ‘Llama’ specifically refers to the Llama 2 version. The ‘XB’ indicates the model’s parameter count, either 7B or 13B. ‘LH’ stands for likelihood, signifying the use of $p_{\theta}(x)$ as the criterion for OOD detection; ‘LR’ denotes likelihood-ratio, referring to the employment of $S(x)$ as outlined in Equation (1) for OOD identification.
@@ 5.1 Far OOD Detection
20NG SST-2 Zhou et al. (2021) 0.978 0.865 0.015
CE (Hendrycks & Gimpel, 2016) 0.981 0.942 0.087
TAPT (Gururangan et al., 2020) 0.981 0.939 0.088
SupCon (Khosla et al., 2020) 0.980 0.943 0.094
Uppaal et al. (2023) 1.000 0.999 0.000
@@
RTE Zhou et al. (2021) 0.956 0.860 0.312
CE (Hendrycks & Gimpel, 2016) 0.945 0.902 0.285
TAPT (Gururangan et al., 2020) 0.919 0.869 0.352
SupCon (Khosla et al., 2020) 0.952 0.914 0.248
Uppaal et al. (2023) 1.000 0.999 0.000
@@
IMDB Zhou et al. (2021) 0.969 0.996 0.144
CE (Hendrycks & Gimpel, 2016) 0.961 0.995 0.206
TAPT (Gururangan et al., 2020) 0.965 0.995 0.159
SupCon (Khosla et al., 2020) 0.970 0.996 0.150
Uppaal et al. (2023) 0.990 0.998 0.012
@@
Table 1: Results of far OOD detection, utilizing the same experimental setup as described by Uppaal et al. (2023). Results for methods not originating from our work are cited directly from Uppaal et al. (2023). This table presents a condensed version; for the complete details, please see Table 5 in Appendix.
5.2 Near OOD Detection
We select the ROSTD (Gangal et al., 2020), SNIPS (Coucke et al., 2018), and CLINC150 (Larson et al., 2019) datasets for our near OOD detection experiments. The ROSTD and CLINC150 datasets are specifically crafted for OOD detection and include designated classes representing OOD data from the same domain. The SNIPS dataset comprises user utterances distributed among seven intent classes, such as GetWeather and RateBook. As it does not inherently provide OOD utterances, we classify the GetWeather and BookRestaurant intents as OOD for the purpose of our experiments. It’s noteworthy that this classification diverges from the one in the study by Jin et al. (2022), which does not explicitly detail their data splitting methodology.
Table 2 demonstrates that the likelihood ratio between the pretrained Llama model and the finetuned Llama model yields the highest performance among unsupervised OOD detection methods. In the case of CLINC150, the supervised OOD detection method introduced by Podolskiy et al. (2021) significantly surpasses our approach, a point that is further discussed in Section 6.

### v2

@@ Abstract
We revisit the likelihood ratio between a pretrained large language model (LLM) and its finetuned variant as a criterion for out-of-distribution (OOD) detection. The intuition behind such a criterion is that, the pretrained LLM has the prior knowledge about OOD data due to its large amount of training data, and once finetuned with the in-distribution data, the LLM has sufficient knowledge to distinguish their difference. Leveraging the power of LLMs, we show that, the likelihood ratio can serve as an effective OOD detection criterion. Moreover, we apply the proposed LLM-based likelihood ratio to detect OOD questions in question-answering (QA) systems, which can be used to improve the performance of specialized LLMs for general questions. Given that likelihood can be easily obtained by the loss functions within contemporary neural network frameworks, it is straightforward to implement this approach in practice. Since both the pretrained LLMs and its various finetuned models are widely available from online platforms such as Hugging Face, our proposed criterion can be effortlessly incorporated for OOD detection without the need for further training. We conduct comprehensive evaluation across on multiple settings, including far OOD, near OOD, spam detection, and QA scenarios, to demonstrate the effectiveness of the method. Code can be found at https://github.com/andiac/LLMOODratio
@@ 1 INTRODUCTION
In OOD detection (Hendrycks and Gimpel,, 2016), the term “in-distribution” refers specifically to the distribution of the training data. In natural language processing, OOD detection has been studied in small and task-specific models for settings such as translation (Xiao et al.,, 2020) and question answering (Lyu et al.,, 2020). With the advancement of large and general models like large language models (LLMs), the scope of training data has significantly broadened, positioning these large models embodied with general knowledge and intelligence as “base models”. In the era of large foundation models (Bommasani et al.,, 2021), the prevalent training paradigm has shifted from end-to-end learning towards finetuning a pretrained base model. This shift calls for a revisiting of the definition of “in-distribution” and its relationship to the training distribution of base models.
There is usually no prior information about the OOD data in conventional OOD detection (Hendrycks and Gimpel,, 2016). Bishop, (1994) introduced the idea that OOD data could be viewed as data coming from a distinct OOD distribution, suggesting the use of the likelihood ratio between this OOD distribution and the in-distribution as a detection criterion. However, given the absence of information about OOD data, Zhang and Wischik, (2022) introduced the concept of OODProxy, a conceptual framework where the OOD distribution is represented by a proxy, incorporating what is known or assumed about the OOD data. From this perspective, it is commonly recognized that many OOD detection methods leveraging likelihood ratios are essentially applying their unique OOD proxy distributions, each reflecting different assumptions or prior knowledge concerning the nature of the OOD data (Ren et al.,, 2019; Serrà et al.,, 2019; Schirrmeister et al.,, 2020; Zhang et al.,, 2021).
In our paper, we revisit [^1] Jin et al., (2022) found this method inferior to contrastive learning with smaller models. Our work shows it outperforms as model scale increases (Section 7). that pretrained base models can function as a repository of prior knowledge for OOD data relative to in-distribution data, effectively acting as OOD proxy distributions. Guided by this insight, we discover that the likelihood ratio between the base model and its finetuned counterpart serves as an effective criterion for detecting OOD data. Moreover, for LLM-based question-answering (QA) systems, the same likelihood ratio excels in detecting OOD questions. By identifying and rejecting these OOD questions, we can greatly enhance the robustness of current QA systems.
The convenience of obtaining likelihood from loss functions in current neural networks enables a simple and straightforward implementation of this method in practice. Moreover, it is worth mentioning that numerous practitioners likely already have both a pretrained and a finetuned LLM at their disposal, with both types of models widely accessible from platforms like Hugging Face. This setup inherently equips them with the capacity for OOD detection without necessitating additional training efforts.
@@ 2 BACKGROUND AND PRELIMINARIES
We start with an in-distribution dataset, denoted as $\mathcal{D}_{\text{in}}$, assuming that the data within $\mathcal{D}_{\text{in}}$ is sampled from an in-distribution probability distribution $p_{\text{in}}$. The objective of OOD detection is to determine whether a given input data $x$ originates from $p_{\text{in}}$. Hendrycks and Gimpel, (2016) were the first to highlight the significance of this problem in deep learning era and introduced a practical benchmark for evaluation. This involves training a model on $\mathcal{D}_{\text{in}}^{\text{train}}$, the training subset of the in-distribution dataset, with the model providing a detection criterion $S$ for identifying OOD data. We then gather a dataset from domains different from $\mathcal{D}_{\text{in}}$, labeled as $\mathcal{D}_{\text{out}}$. The effectiveness of the criterion $S$ is assessed by applying it to data from $\mathcal{D}_{\text{in}}^{\text{test}}\cup\mathcal{D}_{\text{out}}^{\text{test}}$ and evaluating the performance using metrics such as AUROC, AUPR, and FPR95 (Yang et al.,, 2022). These metrics help determine how well $S$ can differentiate between the in-distribution dataset and the OOD dataset. A high performance across these metrics signifies a robust OOD detection capability.
“Supervised” and “Unsupervised” OOD Detection
@@
In Hendrycks and Gimpel, (2016)’s foundational study, it is assumed that for the in-distribution dataset $\mathcal{D}_{\text{in}}$, each data point is accompanied by a classification label. This is why this problem setting is referred to as supervised OOD detection [^2] A possible ambiguity here is that some might think having OOD data is what makes it supervised OOD detection; here we emphasize that “supervised” and “unsupervised” refer to the labels of in-distribution data.. Following Hendrycks’ work, numerous supervised OOD detection methods based on classifiers have been proposed (Lee et al.,, 2018; Liu et al.,, 2020; Wang et al.,, 2022; Sun et al.,, 2022; Zhu et al.,, 2022). In contrast, our work focuses on a different scenario, as we operate under the assumption that no labels are available. This scenario is often referred to as “unsupervised OOD detection” or “OOD detection without in-distribution labels”.
The Paradox in Unsupervised OOD Detection
@@
In the context of unsupervised OOD detection, Nalisnick et al., (2018) revisit Bishop, (1994)’s suggestion that a probabilistic generative model $p_{\theta}$ could model the in-distribution $p_{\text{in}}$, proposing to use the model output $S(x)=p_{\theta}(x)$ as a criterion for evaluating a given input $x$. Surprisingly, Nalisnick et al. discovered that in certain scenario - such as when $\mathcal{D}_{\text{in}}$ is CIFAR10 and $\mathcal{D}_{\text{out}}$ is SVHN, or $\mathcal{D}_{\text{in}}$ is FashionMNIST and $\mathcal{D}_{\text{out}}$ is MNIST — the OOD data received higher $p_{\theta}(x)$ scores than the in-distribution data. This counterintuitive finding has been labeled as a “paradox”.
OOD Proxy
@@
To address the paradox, the studies (Ren et al.,, 2019; Serrà et al.,, 2019; Schirrmeister et al.,, 2020; Zhang et al.,, 2021; Caterini and Loaiza-Ganem,, 2022) have put forward the idea of utilizing the likelihood ratio as the criterion for identifying OOD data. Zhang and Wischik, (2022) integrates these techniques into a comprehensive structure termed the OOD proxy framework. Within this framework, it is posited that in-distribution data can be characterized as samples from a distribution $p_{\text{in}}$, whereas OOD data are samples from a distribution $p_{\text{out}}$. According to the Neyman-Pearson lemma, the likelihood ratio represents the optimal criterion for OOD detection in theory, which is mathematically expressed as:
@@
where obtaining $p_{\text{out}}$ is challenging. To address this, Zhang and Wischik, (2022) introduced the concept of utilizing a proxy distribution, $p_{\text{out}}^{\text{proxy}}$, which incorporates human subjective understanding of the OOD distribution. For instance, Ren et al., (2019) define $p_{\text{out}}^{\text{proxy}}$ as the distribution representing background statistics; Serrà et al., (2019) consider it as the distribution of data compression; Schirrmeister et al., (2020) describe $p_{\text{out}}^{\text{proxy}}$ as a general distribution; and for Zhang et al., (2021), $p_{\text{out}}^{\text{proxy}}$ corresponds to the distribution of a local autoregressive model.
In the OOD proxy framework, the use of likelihood as a detection criterion, following the approach of Bishop, (1994) and Nalisnick et al., (2018), essentially assumes a uniform distribution for $p_{\text{out}}^{\text{proxy}}$. The suboptimal performance associated with this method is not surprising; it highlights the limitations of an improper prior assumption.
@@
Autoregressive language models (Brown et al.,, 2020) are types of probabilistic models that compute the likelihood of a sentence $x=x_{1},\dots,x_{T}$, where $T$ denotes the sentence length and $x_{t}$ represents each word at position $t$. By definition of conditional probability, the likelihood $p(x)=p(x_{1},\dots,x_{T})$ can be calculated as the product of conditional probabilities: $p(x_{1})p(x_{2}|x_{1})p(x_{3}|x_{1},x_{2})\dots p(x_{T}|x_{1},\dots,x_{T-1})$. Each term $p(x_{t}|x_{<t})$ represents the probability of predicting the subsequent word $x_{t}$ given all the previous words $x_{1},\dots,x_{t-1}$. In neural language models, these conditional probabilities are typically modeled using a softmax function over the output vocabulary.
@@
3 PRETRAINED LLM AS A OOD PROXY
@@
This criterion essentially employs the pretrained large language model as an OOD proxy introduced in Section 2. This strategy is particularly practical given the widespread availability of pretrained LLMs. Finetuning these models to adapt their distribution for specific domain contexts is a standard practice, meaning many practitioners may already possess finetuned LLMs that represent the distribution of their specific datasets. With both pretrained and finetuned models at hand, calculating the likelihood ratio becomes straightforward, eliminating the need for additional training. For example, suppose we possess a LLM that has been finetuned on legal documents. Given a new document $x$, we can determine whether it is a legal document by utilizing the likelihood ratio $S(x)$.
@@
Utilizing a pretrained LLM as an OOD proxy stems from the rationale that the OOD proxy distribution should encapsulate general characteristics of OOD data, including prior knowledge or subjective insights about the OOD data. While it’s conceivable to presume OOD data emanate from a uniform distribution devoid of any prior knowledge, evidence presented by Nalisnick et al., (2018) indicates that relying on a uniform distribution as the OOD proxy can be ineffective in practice. In the context of language models, the well-known infinite monkey theorem [^3] The infinite monkey theorem states that a monkey randomly typing on a typewriter for an infinitely long period will almost surely produce any given text, such as the complete works of William Shakespeare. (Borel,, 1913; Eddington,, 2019) suggests that any text could theoretically be generated from a uniform distribution; however, this does not serve as an effective representation of any coherent language. Figure 1 depicts the relationship among sentences within a specific domain, the comprehensive set of human language, and all conceivable character permutations. In fact, coherent human language forms a minor subset of all potential character arrangements. Consequently, the assumption that OOD data ought to represent meaningful human language constitutes a strong prior.
Therefore, we advocate for the use of a pretrained LLM as a more suitable OOD proxy. Given their extensive parameters and training on vast corpora (for example, the Llama-2 model, as mentioned by Touvron et al., (2023), is trained on 7 trillion tokens), it is plausible to consider that an LLM encompasses the breadth of human language. Assuming the LLM estimates the distribution of all human language, it is logical to designate the pretrained LLM as the OOD proxy.
@@ 4 LIKELIHOOD RATIO OOD DETECTION FOR QA SYSTEMS
$S_{q}(q,a)$ $=\frac{p_{\theta}(q)}{p_{\theta^{\prime}}(q)},$
$S_{a}(q,a)$ $=\frac{p_{\theta}(a)}{p_{\theta^{\prime}}(a)},$
$S_{q,a}(q,a)$ $=\frac{p_{\theta}(q,a)}{p_{\theta^{\prime}}(q,a)},$
$S_{a|q}(q,a)$ $=\frac{p_{\theta}(a|q)}{p_{\theta^{\prime}}(a|q)}=\frac{p_{\theta}(q,a)p_{\theta^{\prime}}(q)}{p_{\theta^{\prime}}(q,a)p_{\theta}(q)}$
$=\frac{S_{q,a}(q,a)}{S_{q}(q,a)},$
where $S_{q}$, $S_{a}$, $S_{q,a}$, and $S_{a|q}$ are defined as follows: $S_{q}$ is the likelihood ratio for the question, $S_{a}$ for the answer, $S_{q,a}$ for the question-answer pair, and $S_{a|q}$ for the answer given the question. All these criterion are the ratios between the likelihoods assigned by the finetuned model to those assigned by the base model. It is difficult to determine intuitively which of these four criteria performs best. In the following experimental section, we will empirically demonstrate which criterion is more effective in practice.
Refer to captionRefer to caption
Figure 2: Example question-answer sets produced by MetaMath-7B. The responses to In-D questions are accurate and logical. However, for OOD questions, MetaMath-7B generates unreasonable answers, responding to a straightforward query with unnecessary mathematical calculations or producing repetitive sentences with no useful information. For the complete image, please see Appendix.
@@
20NG SST-2 Zhou et al., (2021) 0.978 0.865 0.015
CE (Hendrycks and Gimpel,, 2016) 0.981 0.942 0.087
TAPT (Gururangan et al.,, 2020) 0.981 0.939 0.088
SupCon (Khosla et al.,, 2020) 0.980 0.943 0.094
Uppaal et al., (2023) 1.000 0.999 0.000
@@
RTE Zhou et al., (2021) 0.956 0.860 0.312
CE (Hendrycks and Gimpel,, 2016) 0.945 0.902 0.285
TAPT (Gururangan et al.,, 2020) 0.919 0.869 0.352
SupCon (Khosla et al.,, 2020) 0.952 0.914 0.248
Uppaal et al., (2023) 1.000 0.999 0.000
@@
IMDB Zhou et al., (2021) 0.969 0.996 0.144
CE (Hendrycks and Gimpel,, 2016) 0.961 0.995 0.206
TAPT (Gururangan et al.,, 2020) 0.965 0.995 0.159
SupCon (Khosla et al.,, 2020) 0.970 0.996 0.150
Uppaal et al., (2023) 0.990 0.998 0.012
@@
Table 1: Results of far OOD detection, utilizing the same experimental setup as described by Uppaal et al., (2023). Results for methods not originating from our work are cited directly from Uppaal et al., (2023). For the complete details, please see Appendix.
5 EXPERIMENTS
We adhere to the definitions of far OOD and near OOD as outlined by Yang et al., (2022) in their work. Near OOD datasets exhibit only a semantic shift from the In-D datasets, whereas far OOD also encompasses a significant covariate (domain) shift. For far OOD evaluations, we designate two distinct datasets as In-D and OOD. For near OOD, we divide a single dataset into two groups: one serving as the In-D with certain classes and the other as OOD with a different set of classes.
Additionally, we demonstrate the capability of our proposed method in detecting OOD instances within the context of spam detection (Labonne and Moran,, 2023) — a practical application for our unsupervised OOD detection technique, especially where In-D labels are absent. We show that our method achieves commendable results in spam detection even without access to any spam data. Moreover, when spam data is available and we further finetune the OOD proxy distribution using this data, our results are competitive with state-of-the-art (SOTA) spam detection algorithms.
Finally, we evaluate our approach within a real question-answering (QA) context, utilizing MetaMath — a Llama-2 model finetuned for math problem-solving, as described by Yu et al., (2023). By implementing the likelihood-ratio-based criteria outlined in Section 4, we find that for specific short questions, having the LLM provide an answer and subsequently applying a criterion that analyzes both the question and answer leads to consistently improved outcomes in identifying OOD questions.
Evaluation Metrics
We employ AUROC (Area Under the Receiver Operating Characteristic curve), AUPR (Area Under the Precision-Recall curve), and FPR95 (False Positive Rate at 95% True Positive Rate) as our evaluation metrics. These metrics are commonly utilized in evaluating the performance of OOD detection methods (Hendrycks and Gimpel,, 2016; Yang et al.,, 2022).